Calcium carbide, CaC2, is an important preliminary chemical for industries producing synthetic fabrics and plastics. CaC2 may be produced by heating calcium oxide with coke: CaO + 3C → CaC2 + CO.


In a reaction, 1.5kg of carbon was reacted with calcium oxide and calcium carbide was produced. calculate the percent yield of calcium carbide.


1
Expert's answer
2021-05-20T07:50:28-0400

Moles of Carbon = 1500g/12g = 125 mol

When 3 mole of C react with 1 mole of CaO

Moles of CaO = 125mol/3 = 41.66 mol

Mass of CaO = 41.66 × 56.077 = 2336.16 = Actual yield

Similarly moles of CaC2 = 41.66 mol

Mass of CaC2 = 41.66 × 64.099 = 2670.36 g = theoretical yield

Percentage yield = "\frac{Actual yield}{theoretical yield} ×100"


Percentage yield = "\frac{2336.16}{2670.36} ×100" = 87.4%